the combined ages of joel his father and grandfather are equal to 100. joel is 25 years younger than his father and 51 years younger than his grandfather. How old are joel, his father and his grandfather?
Found 2 solutions by IWork4Dessert, anomis:
Answer by IWork4Dessert(60)

The first thing to do here is to separate your equation into parts. The first sentence says, "the combined ages of Joel, his father, and his grandfather are equal to 100." This is going to be your equation:
Joel+Father+Grandfather=100
Now you need to find your variables. The problem then says, "Joel is 25 years younger than his father and 51 years younger than his grandfather." Next you break it down.
Joel is 25 years younger than his father: Joel=x-25
Father=x
...and 51 years younger than his grandfather: Now you subtract 25 from 51, which would give you the distance between the ages of the father and the grandfather.
51-25=26
This means that the grandfather is 26 years older than the father. Knowing this, you can have:
Grandfather=x+26
Now you put it all together.
(x-25)+x+(x+26)=100
3x+1=100
3x=99
x=33
Now you know your x, which is the father. Just plug it back in.
x=33
x-25=8
x+26=59
To check your answer, add these numbers.
33+8+59=100
And you're done.

Answer by anomis(18)

J =Joel
J+25 = Father
J+51 = Grandfather
we know that J+F+G=100 => J + J + 25 + J + 51 = 100 => 3J + 76=100 => 3J=100-76
=> 3J = 24 => J = 24/3 = 8
Joel is 8 years old
Joel's father is 8 + 25 = 33y.o
Joel's Grandfather is 8 + 51 = 59y.o
Verification
8 + 33 + 59 = 100